OpenStudy (anonymous):

A vehicle moves in a straight line with an acceleration of 5 km/h2. By how much does the speed change each second? Answer in units of km/h.

OpenStudy (anonymous):

a=(v-u)/t at=v-u. let u=0, and sub t=1.

OpenStudy (anonymous):

so, a=(5-0)/1 a=5 ?

OpenStudy (anonymous):

a=5 (given) \(at=v-u=\Delta v\)

OpenStudy (ksaimouli):

chage in speed= final velocity- initial velocity

OpenStudy (ksaimouli):



OpenStudy (vincent-lyon.fr):

5 km/h² = (5 km/h) / h = (5 km/h) / 3600 s = 5/3600 (km/h)/s \(\Delta v\) = 0.0014 km/h in one second